# Kohn–Sham inversion with mathematical guarantees

### v2

@@ Abstract
We use an exact Moreau-Yosida regularized formulation to obtain the exchange-correlation potential for periodic systems. We reveal a profound connection between rigorous mathematical principles and efficient numerical implementation, which marks the first computation of a Moreau–Yosida-based inversion for physical systems. We develop a mathematically rigorous inversion algorithm which is demonstrated for representative bulk materials, specifically bulk silicon, gallium arsenide, and potassium chloride. This inversion algorithm allows the construction of rigorous error bounds that we are able to verify numerically. This unlocks a new pathway to analyze Kohn–Sham inversion methods, which we expect in turn to foster mathematical approaches for developing approximate functionals.
@@ II.2 The Inversion Formula & Error Bounds
The first term is the error arising from terminating the inversion at $\varepsilon>0$. This term is guaranteed to vanish as $\varepsilon\to 0^{+}$ by [49, Thm. 2] and will be negligible for sufficiently small $\varepsilon$, a behavior also seen in our numerical calculations. Of interest for this article is the second term, sometimes referred to as a density-driven error [68, 69, 70]. Using that $v_{\mathrm{xc}}^{\varepsilon}=\tfrac{1}{\varepsilon}J(\rho_{\mathrm{gs}}^{\varepsilon}-\rho_{\mathrm{gs}})$ and similarly for $\tilde{v}_{\mathrm{xc}}^{\varepsilon}$, the error bound,
@@ III.1 Implementation of the Inversion Algorithm
The uniqueness of the proximal density for a fixed $\varepsilon>0$ implies that $\rho^{\varepsilon}_{\mathrm{gs}}$ can be obtained from $\mathcal{E}$ of Eq. 9 by any minimization procedure of choice. In agreement with the usual KS-DFT implementation for this setting we will employ a parametrization of the proximal density in terms of orthonormal orbitals $\Phi=(\psi_{1},\ldots,\psi_{N_{b}})$, i.e. $\int_{\Omega}\psi_{i}^{\ast}(\mathbf{r})\psi_{j}(\mathbf{r})\mathrm{d}\mathbf{r}=\delta_{ij}$. In this notation we do not make the $k$-points explicit, i.e. the sum over $i$ runs over all orbitals at all reducible k-points used to discretize the Brillouin zone. Moreover in this work we will only consider periodic insulators, such that we can avoid making spin explicit and simply take $N_{b}$ to be the product of the number of electron pairs and the number of $k$-points. In this parametrization the density is thus obtained as
@@
$\mathcal{E}(\Phi,\rho_{\mathrm{gs}})=$ $\,\sum_{i=1}^{N_{b}}\int_{\Omega}\left|\nabla\psi_{i}(\mathbf{r})\right|^{2}\mathrm{d}\mathbf{r}+E_{\mathrm{H}}(\rho_{\Phi})$ (15)
@@ III.4 Noisy inversion and error estimates
By comparing the proximal densities of $\rho_{\mathrm{gs}}$ and that of $\tilde{\rho}_{\mathrm{gs}}$, we can investigate the non-expansiveness of the proximal mapping by computing the ratio $Q_{\varepsilon}(\Delta\rho)$. A plot of $Q_{\varepsilon}$ is shown in LABEL:{fig:ContractionConstantTruncation} for a series of perturbations induced by truncations of the density in the Fourier basis, where the corresponding size of the perturbation $\left\|\Delta\rho\right\|_{\mathcal{D}}$ is shown in Fig. 5. By the proximal map, $Q_{\varepsilon}$ is bounded to the range $[0,1]$. From Fig. 6 we see that for large values of the regularization parameter ($\varepsilon\sim 1$) $Q_{\varepsilon}\ll 1$, and $Q_{\varepsilon}\to 1^{-}$ as $\varepsilon\to 0^{+}$. Additionally, we may obtain estimates for the error bounds on the xc potential, Eqs. 14 and 16, directly from the ratios $R_{\varepsilon}$ and $S_{\varepsilon}$. In particular, Fig. 7 shows the ratios $R_{\varepsilon}$ and $S_{\varepsilon}$ (respectively bottom and top panels) obtained using the same $\tilde{\rho}_{\mathrm{gs}}$ as in Fig. 6. Furthermore, Fig. 7 (top) shows that the ratio $S_{\varepsilon}$ depend on $\varepsilon$ in a very similar manner as $Q_{\varepsilon}$ (see Fig. 6). Note that for large values of $\varepsilon$ the ratio $S_{\varepsilon}$ does not adhere exactly to the bound $S_{\varepsilon}\leq Q_{\varepsilon}$ (Eq. 16), as can be seen by comparing Figs. 6 and 7. However, this is not surprising given that it is obtained from the difference of three quantities that should pointwise be almost zero. For small values of $\varepsilon$, on the other hand, $S_{\varepsilon}$ is in excellent agreement with the bound set by $Q_{\varepsilon}$, also satisfying $S_{\varepsilon}\to 1$ as $\varepsilon\to 0^{+}$. Moreover, Fig. 7 (bottom) shows that the ratio $R_{\varepsilon}$ adheres strictly to the bound set by $Q_{\varepsilon}$ (Eq. 14). In particular, $R_{\varepsilon}$ follow closely the lower bound $R_{\varepsilon}\geq 1-Q_{\varepsilon}$ (Eq. 17). Although obtaining the ratios involved in the error bounds, Eqs. 14, 16 and 17, is not computationally feasible in all practical calculations, our numerical results conform to the theoretical predictions. In fact, Figs. 6 and 7 even indicate that the ratio $Q_{\varepsilon}$ may be estimated by a constant independent of $\Delta\rho$, but parametrically dependent on $\varepsilon$ (and the guiding functional $\mathcal{F}$). Theoretically exploring this observation further marks an extremely promising direction for future research towards more approximate bounds, which could be applied within a practical inversion scheme to make it more efficient and reliable.

### v3

@@ Abstract
We use an exact Moreau-Yosida regularized formulation to obtain the exchange-correlation potential for periodic systems. We reveal a profound connection between rigorous mathematical principles and efficient numerical implementation, which marks the first computation of a Moreau–Yosida-based inversion for physical systems. We develop a mathematically rigorous inversion algorithm which is demonstrated for representative bulk materials, specifically bulk silicon, gallium arsenide, and potassium chloride. Our inversion algorithm allows the construction of rigorous error bounds that we are able to verify numerically. This unlocks a new pathway to analyze Kohn–Sham inversion methods, which we expect in turn to foster mathematical approaches for developing approximate functionals.
@@ II.2 The Inversion Formula & Error Bounds
The first term is the error arising from terminating the inversion at $\varepsilon>0$. This term is guaranteed to vanish as $\varepsilon\to 0^{+}$ by [49, Thm. 2] and will be negligible for sufficiently small $\varepsilon$, a behavior also seen in our numerical calculations. Of interest in this article is the second term in Eq. 13, i.e., the error arising from the use of an inexact density. In the literature, errors in the energy caused by the use of perturbed densities are referred to as density-driven errors [68, 69, 70]. We will therefore borrow the term “density-driven” when we speak of errors in the potentials caused by inexact densities. Using that $v_{\mathrm{xc}}^{\varepsilon}=\tfrac{1}{\varepsilon}J(\rho_{\mathrm{gs}}^{\varepsilon}-\rho_{\mathrm{gs}})$ and similarly for $\tilde{v}_{\mathrm{xc}}^{\varepsilon}$, the error bound,
@@ III.1 Implementation of the Inversion Algorithm
The uniqueness of the proximal density for a fixed $\varepsilon>0$ implies that $\rho^{\varepsilon}_{\mathrm{gs}}$ can be obtained from $\mathcal{E}$ of Eq. 9 by any minimization procedure of choice. In agreement with the usual KS-DFT implementation for this setting we will employ a parametrization of the proximal density in terms of orthonormal orbitals $\Phi=(\psi_{1},\ldots,\psi_{N_{b}})$, i.e. $\int_{\Omega}\psi_{i}^{\ast}\psi_{j}=\delta_{ij}$. In this notation we do not make the $k$-points explicit, i.e. the sum over $i$ runs over all orbitals at all reducible k-points used to discretize the Brillouin zone. Moreover in this work we will only consider periodic insulators, such that we can avoid making spin explicit and simply take $N_{b}$ to be the product of the number of electron pairs and the number of $k$-points. In this parametrization the density is thus obtained as
@@
$\mathcal{E}(\Phi,\rho_{\mathrm{gs}})=$ $\,\sum_{i=1}^{N_{b}}\int_{\Omega}\left|\nabla\psi_{i}\right|^{2}+E_{\mathrm{H}}(\rho_{\Phi})$ (15)
@@ III.4 Noisy inversion and error estimates
By comparing the proximal densities of $\rho_{\mathrm{gs}}$ and that of $\tilde{\rho}_{\mathrm{gs}}$, we can investigate the non-expansiveness of the proximal mapping by computing the ratio $Q_{\varepsilon}(\Delta\rho)$. A plot of $Q_{\varepsilon}$ is shown in LABEL:{fig:ContractionConstantTruncation} for a series of perturbations induced by truncations of the density in the Fourier basis, where the corresponding size of the perturbation $\left\|\Delta\rho\right\|_{\mathcal{D}}$ is shown in Fig. 5. By the proximal map, $Q_{\varepsilon}$ is bounded to the range $[0,1]$. From Fig. 6 we see that for large values of the regularization parameter ($\varepsilon\sim 1$) $Q_{\varepsilon}\ll 1$, and $Q_{\varepsilon}\to 1^{-}$ as $\varepsilon\to 0^{+}$. Additionally, we may obtain estimates for the error bounds on the xc potential, Eqs. 14 and 16, directly from the ratios $R_{\varepsilon}$ and $S_{\varepsilon}$. In particular, Fig. 7 shows the ratios $R_{\varepsilon}$ and $S_{\varepsilon}$ (respectively bottom and top panels) obtained using the same $\tilde{\rho}_{\mathrm{gs}}$ as in Fig. 6. Furthermore, Fig. 7 (top) shows that the ratio $S_{\varepsilon}$ depend on $\varepsilon$ in a very similar manner as $Q_{\varepsilon}$ (see Fig. 6). Note that for large values of $\varepsilon$ the ratio $S_{\varepsilon}$ does not adhere exactly to the bound $S_{\varepsilon}\leq Q_{\varepsilon}$ (Eq. 16), as can be seen by comparing Figs. 6 and 7. However, this is not surprising given that it is obtained from the difference of three quantities that should pointwise be almost zero. For small values of $\varepsilon$, on the other hand, $S_{\varepsilon}$ is in excellent agreement with the bound set by $Q_{\varepsilon}$, also satisfying $S_{\varepsilon}\to 1$ as $\varepsilon\to 0^{+}$. Moreover, Fig. 7 (bottom) shows that the ratio $R_{\varepsilon}$ adheres strictly to the bound set by $Q_{\varepsilon}$ (Eq. 14). In particular, $R_{\varepsilon}$ follow closely the lower bound $R_{\varepsilon}\geq 1-Q_{\varepsilon}$ (Eq. 17). Although obtaining the ratios involved in the error bounds, Eqs. 14, 16 and 17, is not computationally feasible in all practical calculations, our numerical results conform to the theoretical predictions. In fact, Figs. 6 and 7 even indicate that the ratio $Q_{\varepsilon}$ may be estimated by a constant independent of $\Delta\rho$, but parametrically dependent on $\varepsilon$ (and the guiding functional $\mathcal{F}$). Theoretically exploring this observation further marks a promising direction for future research towards more approximate bounds, which could be applied within a practical inversion scheme to make it more efficient and reliable. It is expected that improvements of the error bounds will enhance their applicability in practical calculations. Continued investigations will reveal if these first, but stringent, error estimates can be made useful in a broader sense.